# Graph-based Square-Root Estimation for Sparse Linear Regression

### v1

@@ 4.1 Simulation data
Table 8: Performance comparison of estimation, prediction and model selection for Example 3 with uniform distribution noise.
$L_{2}$ distance RPE FPR FNR
(I) 20/20/400 Ridge 9.235 (0.635) 7.242 (1.596) 0.965 (0.014) 0.004 (0.016)
Lasso 11.223 (1.375) 6.974 (2.660) 0.095 (0.022) 0.528 (0.125)
Alasso 11.030 (1.505) 6.383 (2.996) 0.091 (0.023) 0.487 (0.136)
Enet 8.867 (0.967) 6.761 (1.906) 0.410 (0.052) 0.107 (0.097)
SRL 10.934 (1.337) 6.374 (2.637) 0.168 (0.030) 0.420 (0.160)
SRIG-o 10.386 (0.339) 11.342 (1.038) 0.845 (0.070) 0.021 (0.044)
SRIG 9.817 (0.478) 9.938 (1.246) 0.908 (0.050) 0.015 (0.036)
GSRE-o 7.045 (1.627) 4.148 (2.630) 0.271 (0.109) 0.056 (0.105)
GSRE 6.253 (1.542) 3.297 (2.353) 0.298 (0.119) 0.020 (0.074)
(II) 40/40/400 Ridge 7.613 (0.545) 3.755 (1.062) 0.960 (0.012) 0.000 (0.000)
Lasso 7.582 (0.884) 2.775 (1.029) 0.054 (0.022) 0.187 (0.086)
Alasso 7.546 (0.995) 2.077 (0.830) 0.039 (0.015) 0.172 (0.089)
Enet 6.924 (0.839) 4.381 (1.315) 0.314 (0.054) 0.017 (0.035)
SRL 7.334 (0.972) 2.562 (1.028) 0.061 (0.023) 0.153 (0.092)
SRIG-o 10.282 (0.284) 11.274 (1.056) 0.508 (0.086) 0.012 (0.029)
SRIG 9.594 (0.490) 9.733 (1.305) 0.602 (0.111) 0.000 (0.000)
GSRE-o 5.469 (1.250) 2.684 (1.719) 0.035 (0.027) 0.012 (0.032)
GSRE 4.490 (0.984) 1.628 (0.959) 0.034 (0.030) 0.000 (0.000)
(III) 60/60/400 Ridge 6.717 (0.618) 2.238 (0.460) 0.961 (0.014) 0.000 (0.000)
Lasso 5.610 (0.919) 1.487 (0.556) 0.047 (0.026) 0.055 (0.061)
Alasso 6.034 (0.989) 1.178 (0.439) 0.024 (0.012) 0.084 (0.064)
Enet 6.684 (0.841) 4.430 (1.261) 0.193 (0.054) 0.004 (0.016)
SRL 5.584 (0.906) 1.465 (0.506) 0.046 (0.025) 0.052 (0.059)
SRIG-o 4.912 (0.815) 2.170 (0.871) 0.024 (0.012) 0.000 (0.000)
SRIG 4.010 (0.606) 1.268 (0.515) 0.044 (0.010) 0.000 (0.000)
GSRE-o 3.899 (0.669) 1.120 (0.514) 0.029 (0.021) 0.000 (0.000)
GSRE 3.642 (0.692) 0.956 (0.514) 0.028 (0.016) 0.000 (0.000)
In summary, the simulation results indicate that methods incorporating graphical structure among predictors outperform popular penalization methods. In various noise environments, the proposed GSRE method, which combines square-root estimation with graph penalty, demonstrates superior performance in estimation, prediction and model selection. This result highlights the general applicability and reliability of the GSRE method in statistical modeling.
@@ 4.2 Real data
The first set is the bodyfat2 data, which originates from the bodyfat data in the LIBSVM database. This data is suitable for large-scale regression problems and can be available at https://www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets. For this study, we first preprocess the data by removing zero columns. Additionally, we use the method of [22] to expand the data by applying polynomial basis functions to the original features. The final digit 2 in bodyfat2 indicates that we use second-order polynomials to generate the basis functions. After expansion, we obtain the bodyfat2 data with a sample size of 252 and a dimensionality of 120. To satisfy the high-dimensional assumption, i.e., $p>n$, we randomly select 100 samples. Therefore the size of the data set we employ is $100\times 120$.
@@
The second set is the miRNA data, which is a small noncoding RNAs that play a crucial role in various biological processes such as development, cellular differentiation, cell cycle regulation and programmed cell death. Recent studies have shown a strong association between miRNA and cancer [21, 25, 19]. Here, we use the miRNA expression data for thymic cancer, which can be downloaded from R package. This data is sourced from thymoma tissue samples in the Cancer Genome Atlas database. This data set provides miRNA expression values for 113 patients across 1,881 miRNAs. After removing duplicate features and filtering out low-expression genes, we obtained expression values for 136 miRNAs. Consequently, the data size is $113\times 136$, which satisfies the high-dimensional assumption of $p>n$. We also consider survival time as the response variable $y$.
Figure 2: Estimated graphs of real data. Left:bodyfat2. Right:miRNA.
@@
Before inputting the preprocessed data into the algorithm for iteration, we perform normalization on the data sets. We then use the graphical lasso [18] to estimate the graphical structure of the predictors, as detailed in Figure 2. It is evident that there is a high degree of correlation among the predictors in both data sets.
@@
In the specific experiments, we use ten-fold cross-validation to evaluate different methods. Here, we randomly select nine folds of data as the training set and use the remaining data as the test set. All models are fitted on the training set and evaluated on the test data. To select the tuning parameters for different methods, we use five-fold cross-validation on the training set. We repeat the experiment 10 times to avoid bias introduced by random partitioning. Here, we uniformly use mean squared error (MSE) to measure prediction accuracy.
Figure 3: Comparison of MSE for various methods on real data. Left:bodyfat2. Right:miRNA.
@@
Figure 3 shows box plots of MSE from 10 repeated experiments for different methods on two real data sets. Figure 3 (a) illustrates the performance of various methods on the bodyfa2 data. The results indicate that GSRE significantly outperforms Ridge, Lasso, Alasso, Enet and SRIG in terms of MSE scores. Although the box sizes of GSRE and SRL are similar, GSRE’s median is noticeably lower than that of SRL. Figure 3 (b) presents the numerical performance of different methods on the miRNA data. This figure clearly shows that GSRE’s median is significantly lower than that of other methods. These results indicate that the GSRE method has superior numerical performance compared to other methods and is an ideal choice for achieving accurate predictions.

### v2

@@ 4.1 Simulation data
[h] Performance comparison of estimation, prediction and model selection for Example 1&2 with T-distribution noise. Example 1 Example 2 $L_{2}$ distance RPE FPR FNR MCC $L_{2}$ distance RPE FPR FNR MCC (I) 20/20/400 Ridge 6.148(1.203) 2.124(2.710) 0.948(0.019) 0.000(0.000) 0.089(0.017) 6.104(1.347) 2.683(2.486) 0.953(0.016) 0.001(0.009) 0.082(0.025) Lasso 10.737(1.884) 2.528(2.470) 0.039(0.022) 0.405(0.144) 0.606(0.153) 9.997(1.916) 4.221(2.683) 0.058(0.027) 0.408(0.146) 0.553(0.160) Alasso 10.612(2.038) 1.937(2.471) 0.026(0.024) 0.373(0.142) 0.669(0.160) 9.908(2.205) 4.109(2.916) 0.036(0.023) 0.403(0.165) 0.614(0.162) Enet 7.129(2.365) 1.344(1.761) 0.116(0.036) 0.192(0.136) 0.599(0.144) 6.830(1.950) 1.922(1.587) 0.113(0.030) 0.159(0.111) 0.627(0.118) SRL 10.439(1.944) 2.440(2.333) 0.108(0.024) 0.337(0.147) 0.504(0.136) 10.056(1.668) 4.310(2.286) 0.113(0.026) 0.344(0.136) 0.490(0.142) SRIG-o 8.822(0.654) 13.761(2.319) 0.559(0.062) 0.000(0.000) 0.326(0.037) 9.213(0.500) 12.015(1.639) 0.550(0.054) 0.005(0.030) 0.328(0.034) SRIG 8.822(0.654) 13.761(2.319) 0.559(0.062) 0.000(0.000) 0.326(0.037) 9.211(0.504) 12.010(1.645) 0.550(0.054) 0.005(0.030) 0.328(0.034) GSRE-o 2.280(0.762) 0.450(0.835) 0.124(0.037) 0.000(0.000) 0.722(0.059) 3.127(1.211) 0.936(1.616) 0.156(0.053) 0.005(0.038) 0.672(0.075) GSRE 2.280(0.762) 0.450(0.835) 0.124(0.037) 0.000(0.000) 0.722(0.059) 3.127(1.211) 0.936(1.616) 0.156(0.053) 0.005(0.038) 0.672(0.075) (II) 40/40/400 Ridge 4.991(1.161) 1.443(1.676) 0.962(0.019) 0.000(0.000) 0.075(0.020) 5.440(0.887) 1.856(0.849) 0.969(0.016) 0.000(0.000) 0.067(0.019) Lasso 6.917(2.493) 1.288(1.422) 0.020(0.022) 0.168(0.167) 0.833(0.126) 5.360(2.606) 1.553(1.825) 0.024(0.019) 0.119(0.151) 0.854(0.101) Alasso 6.379(2.773) 0.591(1.225) 0.011(0.027) 0.121(0.142) 0.894(0.149) 4.424(2.505) 0.577(0.669) 0.010(0.012) 0.068(0.107) 0.925(0.081) Enet 3.667(1.299) 1.713(1.899) 0.110(0.054) 0.001(0.009) 0.750(0.088) 4.323(1.399) 1.930(1.813) 0.119(0.053) 0.006(0.019) 0.732(0.084) SRL 6.838(2.383) 1.159(1.202) 0.021(0.022) 0.148(0.140) 0.843(0.121) 5.249(2.256) 1.261(1.102) 0.028(0.020) 0.097(0.121) 0.857(0.083) SRIG-o 2.661(1.171) 1.345(2.103) 0.008(0.039) 0.000(0.000) 0.983(0.074) 3.022(0.827) 1.163(0.882) 0.005(0.021) 0.000(0.000) 0.987(0.049) SRIG 2.661(1.171) 1.345(2.103) 0.008(0.039) 0.000(0.000) 0.983(0.074) 3.020(0.829) 1.160(0.884) 0.005(0.021) 0.000(0.000) 0.987(0.049) GSRE-o 1.749(0.874) 0.382(0.997) 0.009(0.009) 0.000(0.000) 0.972(0.028) 1.872(0.723) 0.280(0.479) 0.015(0.011) 0.000(0.000) 0.954(0.032) GSRE 1.749(0.874) 0.382(0.997) 0.009(0.009) 0.000(0.000) 0.972(0.028) 1.872(0.723) 0.280(0.479) 0.015(0.011) 0.000(0.000) 0.954(0.032) (III) 60/60/400 Ridge 4.661(0.562) 1.193(0.437) 0.954(0.017) 0.000(0.000) 0.083(0.018) 5.243(0.981) 1.642(0.973) 0.954(0.016) 0.000(0.000) 0.084(0.016) Lasso 4.041(1.313) 0.279(0.190) 0.016(0.015) 0.015(0.037) 0.944(0.048) 3.524(2.231) 0.758(1.529) 0.018(0.021) 0.043(0.130) 0.920(0.112) Alasso 3.966(2.020) 0.139(0.129) 0.003(0.006) 0.045(0.080) 0.964(0.051) 3.485(3.752) 0.658(2.436) 0.016(0.064) 0.031(0.084) 0.950(0.156) Enet 2.294(0.455) 0.476(0.247) 0.081(0.040) 0.000(0.000) 0.802(0.078) 3.170(1.341) 1.105(1.343) 0.077(0.054) 0.007(0.034) 0.809(0.106) SRL 3.801(1.291) 0.251(0.184) 0.022(0.017) 0.011(0.031) 0.929(0.051) 3.335(2.029) 0.552(0.904) 0.029(0.061) 0.024(0.085) 0.913(0.150) SRIG-o 1.815(0.333) 0.483(0.197) 0.001(0.004) 0.000(0.000) 0.996(0.011) 2.247(0.744) 0.614(0.587) 0.005(0.023) 0.000(0.000) 0.989(0.054) SRIG 1.815(0.333) 0.483(0.197) 0.001(0.004) 0.000(0.000) 0.996(0.011) 2.250(0.742) 0.615(0.587) 0.005(0.023) 0.000(0.000) 0.989(0.054) GSRE-o 1.331(0.297) 0.091(0.070) 0.006(0.010) 0.000(0.000) 0.980(0.030) 1.591(0.688) 0.194(0.330) 0.008(0.021) 0.000(0.000) 0.979(0.050) GSRE 1.331(0.297) 0.091(0.070) 0.006(0.010) 0.000(0.000) 0.980(0.030) 1.591(0.688) 0.194(0.330) 0.008(0.021) 0.000(0.000) 0.979(0.050)
{sidewaystable}
[h] Performance comparison of estimation, prediction and model selection for Example 1&2 with Laplace noise. Example 1 Example 2 $L_{2}$ distance RPE FPR FNR MCC $L_{2}$ distance RPE FPR FNR MCC (I) 20/20/400 Ridge 6.584(1.155) 2.749(2.811) 0.947(0.021) 0.000(0.000) 0.089(0.019) 6.826(1.336) 3.229(2.910) 0.952(0.017) 0.001(0.009) 0.083(0.022) Lasso 12.095(1.424) 3.178(2.209) 0.059(0.028) 0.487(0.097) 0.491(0.133) 11.222(1.616) 4.903(3.013) 0.071(0.019) 0.445(0.123) 0.491(0.125) Alasso 12.167(1.531) 2.970(3.385) 0.037(0.023) 0.464(0.109) 0.567(0.135) 10.992(1.709) 4.591(2.836) 0.051(0.020) 0.431(0.137) 0.551(0.140) Enet 9.386(1.700) 2.662(2.712) 0.148(0.033) 0.285(0.116) 0.481(0.118) 8.474(1.316) 2.774(1.608) 0.140(0.030) 0.211(0.106) 0.546(0.108) SRL 12.388(1.566) 3.751(3.120) 0.125(0.024) 0.451(0.116) 0.387(0.115) 11.617(1.812) 5.419(2.589) 0.128(0.022) 0.429(0.125) 0.399(0.126) SRIG-o 8.867(0.652) 13.907(2.361) 0.568(0.069) 0.007(0.047) 0.316(0.048) 9.196(0.560) 11.974(1.763) 0.568(0.058) 0.004(0.021) 0.318(0.033) SRIG 8.867(0.652) 13.907(2.361) 0.568(0.069) 0.007(0.047) 0.316(0.048) 9.194(0.565) 11.968(1.772) 0.568(0.058) 0.004(0.021) 0.318(0.033) GSRE-o 3.022(0.890) 1.145(1.178) 0.129(0.035) 0.000(0.000) 0.713(0.058) 4.029(1.139) 1.518(1.721) 0.153(0.047) 0.007(0.047) 0.675(0.069) GSRE 3.022(0.890) 1.145(1.178) 0.129(0.035) 0.000(0.000) 0.713(0.058) 4.029(1.139) 1.518(1.721) 0.153(0.047) 0.007(0.047) 0.675(0.069) (II) 40/40/400 Ridge 6.059(0.749) 2.098(1.317) 0.965(0.021) 0.000(0.000) 0.072(0.020) 6.953(0.855) 2.900(1.155) 0.964(0.016) 0.001(0.009) 0.071(0.023) Lasso 9.571(1.366) 2.585(1.339) 0.014(0.013) 0.333(0.104) 0.741(0.098) 8.235(1.267) 3.047(1.656) 0.023(0.018) 0.261(0.116) 0.761(0.102) Alasso 9.571(1.582) 0.988(0.541) 0.016(0.016) 0.239(0.100) 0.799(0.091) 8.483(1.328) 1.763(0.912) 0.029(0.019) 0.209(0.097) 0.779(0.096) Enet 4.760(0.970) 3.227(1.599) 0.100(0.035) 0.003(0.013) 0.761(0.063) 5.722(1.018) 3.432(1.689) 0.126(0.034) 0.019(0.036) 0.705(0.065) SRL 9.297(1.569) 2.495(1.577) 0.018(0.019) 0.304(0.131) 0.748(0.111) 8.059(1.234) 3.106(1.921) 0.023(0.018) 0.253(0.128) 0.765(0.112) SRIG-o 3.140(1.194) 1.783(1.760) 0.024(0.052) 0.000(0.000) 0.943(0.099) 4.341(1.509) 2.696(2.218) 0.076(0.088) 0.000(0.000) 0.839(0.161) SRIG 3.140(1.194) 1.783(1.760) 0.024(0.052) 0.000(0.000) 0.943(0.099) 4.338(1.503) 2.687(2.201) 0.076(0.088) 0.000(0.000) 0.840(0.162) GSRE-o 2.380(0.575) 0.702(0.593) 0.011(0.011) 0.000(0.000) 0.966(0.033) 3.046(0.727) 0.834(0.634) 0.014(0.016) 0.000(0.000) 0.957(0.044) GSRE 2.380(0.575) 0.702(0.593) 0.011(0.011) 0.000(0.000) 0.966(0.033) 3.051(0.730) 0.833(0.633) 0.014(0.016) 0.000(0.000) 0.957(0.044) (III) 60/60/400 Ridge 6.078(0.724) 2.109(0.569) 0.958(0.014) 0.000(0.000) 0.080(0.015) 6.470(0.743) 2.411(0.783) 0.958(0.012) 0.000(0.000) 0.080(0.012) Lasso 7.711(1.308) 1.454(0.721) 0.007(0.011) 0.196(0.118) 0.858(0.078) 6.464(1.403) 1.901(1.465) 0.010(0.011) 0.129(0.127) 0.889(0.077) Alasso 8.876(1.556) 0.662(0.279) 0.013(0.015) 0.224(0.089) 0.818(0.089) 7.267(1.357) 1.110(0.504) 0.021(0.017) 0.153(0.082) 0.840(0.081) Enet 3.858(0.585) 2.108(0.805) 0.052(0.025) 0.000(0.000) 0.861(0.058) 4.840(1.090) 2.678(1.641) 0.047(0.022) 0.008(0.029) 0.866(0.055) SRL 7.476(1.269) 1.212(0.600) 0.012(0.012) 0.153(0.098) 0.866(0.074) 6.220(1.298) 1.642(0.958) 0.015(0.013) 0.093(0.085) 0.895(0.067) SRIG-o 2.295(0.414) 0.767(0.334) 0.002(0.005) 0.000(0.000) 0.994(0.016) 2.789(0.562) 0.829(0.474) 0.003(0.006) 0.000(0.000) 0.991(0.019) SRIG 2.295(0.414) 0.767(0.334) 0.002(0.005) 0.000(0.000) 0.994(0.016) 2.794(0.553) 0.829(0.474) 0.003(0.006) 0.000(0.000) 0.991(0.019) GSRE-o 2.055(0.331) 0.453(0.226) 0.004(0.007) 0.000(0.000) 0.988(0.021) 2.543(0.424) 0.518(0.269) 0.006(0.009) 0.000(0.000) 0.980(0.028) GSRE 2.055(0.331) 0.453(0.226) 0.004(0.007) 0.000(0.000) 0.988(0.021) 2.547(0.418) 0.518(0.269) 0.006(0.009) 0.000(0.000) 0.980(0.028)
{sidewaystable}
[h] Performance comparison of estimation, prediction and model selection for Example 1&2 with uniform noise. Example 1 Example 2 $L_{2}$ distance RPE FPR FNR MCC $L_{2}$ distance RPE FPR FNR MCC (I) 20/20/400 Ridge 6.636(1.193) 2.672(2.646) 0.952(0.016) 0.003(0.013) 0.081(0.029) 6.724(1.175) 2.989(2.248) 0.955(0.014) 0.001(0.009) 0.080(0.028) Lasso 12.319(1.584) 3.090(2.225) 0.064(0.032) 0.477(0.110) 0.490(0.141) 11.318(1.622) 5.450(3.316) 0.067(0.025) 0.467(0.133) 0.485(0.148) Alasso 12.449(1.728) 3.101(3.412) 0.044(0.024) 0.476(0.110) 0.537(0.136) 11.236(1.514) 4.815(2.851) 0.053(0.021) 0.448(0.136) 0.533(0.147) Enet 9.433(1.293) 2.506(2.153) 0.151(0.027) 0.284(0.102) 0.477(0.103) 8.590(1.360) 2.847(2.307) 0.139(0.027) 0.216(0.096) 0.544(0.100) SRL 12.583(1.426) 3.788(3.163) 0.128(0.020) 0.452(0.123) 0.381(0.119) 11.673(1.773) 5.453(2.772) 0.128(0.028) 0.435(0.138) 0.395(0.140) SRIG-o 8.866(0.659) 13.900(2.364) 0.572(0.071) 0.007(0.047) 0.314(0.050) 9.216(0.540) 12.024(1.726) 0.558(0.061) 0.004(0.021) 0.324(0.034) SRIG 8.866(0.659) 13.900(2.364) 0.572(0.071) 0.007(0.047) 0.314(0.050) 9.215(0.542) 12.022(1.730) 0.558(0.061) 0.004(0.021) 0.324(0.034) GSRE-o 3.198(1.048) 1.371(1.624) 0.138(0.042) 0.000(0.000) 0.701(0.065) 4.123(1.163) 1.564(1.840) 0.144(0.040) 0.007(0.047) 0.687(0.062) GSRE 3.198(1.048) 1.371(1.624) 0.138(0.042) 0.000(0.000) 0.701(0.065) 4.124(1.162) 1.564(1.840) 0.144(0.040) 0.007(0.047) 0.687(0.062) (II) 40/40/400 Ridge 6.257(0.727) 2.131(1.253) 0.965(0.016) 0.000(0.000) 0.072(0.018) 6.886(0.836) 2.852(1.260) 0.964(0.017) 0.001(0.009) 0.069(0.030) Lasso 9.781(1.313) 2.646(1.412) 0.016(0.015) 0.352(0.105) 0.720(0.097) 8.167(1.232) 2.913(1.622) 0.024(0.015) 0.239(0.108) 0.773(0.096) Alasso 9.858(1.525) 1.022(0.461) 0.021(0.013) 0.253(0.103) 0.770(0.091) 8.351(1.316) 1.634(0.793) 0.028(0.014) 0.205(0.088) 0.784(0.076) Enet 4.743(0.911) 3.097(1.494) 0.109(0.036) 0.003(0.013) 0.746(0.063) 5.658(0.952) 3.311(1.576) 0.127(0.031) 0.017(0.040) 0.704(0.059) SRL 9.483(1.613) 2.447(1.423) 0.020(0.017) 0.337(0.127) 0.717(0.104) 8.017(1.287) 3.007(1.883) 0.023(0.015) 0.237(0.119) 0.775(0.092) SRIG-o 3.270(1.305) 1.989(2.091) 0.034(0.065) 0.000(0.000) 0.924(0.120) 4.196(1.332) 2.438(1.887) 0.070(0.089) 0.000(0.000) 0.854(0.160) SRIG 3.270(1.305) 1.989(2.091) 0.034(0.065) 0.000(0.000) 0.924(0.120) 4.192(1.330) 2.426(1.889) 0.069(0.090) 0.000(0.000) 0.856(0.161) GSRE-o 2.446(0.584) 0.736(0.602) 0.012(0.012) 0.000(0.000) 0.965(0.036) 2.945(0.579) 0.724(0.481) 0.016(0.015) 0.000(0.000) 0.953(0.042) GSRE 2.446(0.584) 0.736(0.602) 0.012(0.012) 0.000(0.000) 0.965(0.036) 2.950(0.584) 0.724(0.480) 0.016(0.015) 0.000(0.000) 0.953(0.042) (III) 60/60/400 Ridge 6.174(0.742) 2.215(0.538) 0.957(0.016) 0.000(0.000) 0.080(0.015) 6.426(0.720) 2.424(0.832) 0.959(0.011) 0.000(0.000) 0.079(0.012) Lasso 8.016(1.343) 1.559(0.765) 0.008(0.010) 0.229(0.114) 0.830(0.074) 6.463(1.353) 1.917(1.353) 0.010(0.013) 0.140(0.126) 0.883(0.077) Alasso 9.167(1.393) 0.698(0.263) 0.011(0.012) 0.243(0.079) 0.814(0.078) 7.247(1.284) 1.080(0.461) 0.017(0.014) 0.157(0.085) 0.848(0.078) Enet 3.956(0.664) 2.250(0.949) 0.052(0.025) 0.000(0.000) 0.859(0.057) 4.776(1.123) 2.602(1.686) 0.049(0.027) 0.009(0.036) 0.862(0.068) SRL 7.647(1.211) 1.252(0.551) 0.013(0.013) 0.164(0.098) 0.858(0.065) 6.158(1.142) 1.598(0.878) 0.016(0.016) 0.089(0.089) 0.897(0.060) SRIG-o 2.341(0.423) 0.789(0.357) 0.002(0.004) 0.000(0.000) 0.994(0.014) 2.796(0.590) 0.846(0.516) 0.002(0.005) 0.000(0.000) 0.993(0.016) SRIG 2.341(0.423) 0.789(0.357) 0.002(0.004) 0.000(0.000) 0.994(0.014) 2.802(0.584) 0.846(0.516) 0.002(0.005) 0.000(0.000) 0.993(0.016) GSRE-o 2.112(0.340) 0.463(0.229) 0.005(0.008) 0.000(0.000) 0.985(0.023) 2.516(0.453) 0.507(0.312) 0.006(0.010) 0.000(0.000) 0.983(0.030) GSRE 2.112(0.340) 0.463(0.229) 0.005(0.008) 0.000(0.000) 0.985(0.023) 2.521(0.450) 0.508(0.312) 0.006(0.010) 0.000(0.000) 0.983(0.030)
@@ 4.2 Real data
The first set is the bodyfat2 data, which originates from the bodyfat data in the LIBSVM database. This data is suitable for large-scale regression problems and can be available at https://www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets. For this study, we first preprocess the data by removing zero columns. Additionally, we use the method of Huang et al. (2010) to expand the data by applying polynomial basis functions to the original features. The final digit 2 in bodyfat2 indicates that we use second-order polynomials to generate the basis functions. After expansion, we obtain the bodyfat2 data with a sample size of 252 and a dimensionality of 120. To satisfy the high-dimensional assumption, i.e., $p>n$, we randomly select 100 samples. Therefore, the size of the data set we employ is 100×120100$\times$ 120100 × 120.
@@
The second set is the miRNA data, which is a small noncoding RNAs that play a crucial role in various biological processes such as development, cellular differentiation, cell cycle regulation and programmed cell death. Recent studies have shown a strong association between miRNA and cancer (Hayes et al., 2014; Kong et al., 2012; Garzon et al., 2009). Here, we use the miRNA expression data for thymic cancer, which can be downloaded from R package. This data is sourced from thymoma tissue samples in the Cancer Genome Atlas database. This data set provides miRNA expression values for 113 patients across 1,881 miRNAs. After removing duplicate features and filtering out low-expression genes, we obtain expression values for 136 miRNAs. Consequently, the data size is 113×136113$\times$ 136113 × 136, which satisfies the high-dimensional assumption of $p>n$. We also consider survival time as the response variable $y$.
(a) bodyfat2
(b) miRNA
Figure 2: Estimated graphs of real data.
@@
Before inputting the preprocessed data into the algorithm for iteration, we perform normalization on the data sets. We then use the graphical lasso (Friedman et al., 2008) to estimate the graphical structure of the predictors, as detailed in Figure 2. It is evident that there is a high degree of correlation among the predictors in both data sets.
@@
In the specific experiments, we use ten-fold cross-validation to evaluate different methods. Here, we randomly select nine folds of data as the training set and use the remaining data as the test set. All models are fitted on the training set and evaluated on the test data. To select the tuning parameters for different methods, we use five-fold cross-validation on the training set. The experiment was repeated 10 times to avoid biases caused by random partitioning. Here, we uniformly use mean squared error (MSE) to measure prediction accuracy. Figure 3 shows box plots of MSE for different methods on two real data sets.
(a) bodyfat2
(b) miRNA
Figure 3: Comparison of MSE for various methods on real data.
@@
Figure 3(a) illustrates the performance of various methods on the bodyfa2 data. The results indicate that GSRE significantly outperforms the other six methods in terms of MSE scores. Figure 3(b) presents the numerical performance of different methods on the miRNA data. This figure clearly shows that the performance of GSRE is significantly superior to that of Ridge, Lasso, Alasso, Enet and SRIG. Although the boxes of GSRE and SRL are both at the bottom, the median of GSRE is lower than that of SRL. These results indicate that the GSRE method has superior numerical performance compared to other methods and is an ideal choice for achieving accurate predictions.